OpenStudy (anonymous):

Write a single rule for a combination of these two translations: (x, y) to (x - 1, y + 3) (x, y) to (x + 4, y - 1)

OpenStudy (ash2326):

First translation (x, y) to (x - 1, y + 3) Then (x, y) to (x + 4, y - 1) Let's combine, for the second translation x=x-1 and y=y+3 so ((x-1)+4, (y+3)-1)) Let's simplify (x+3, y+2) so this is the common translation. Do you get this?

OpenStudy (ash2326):

@Megan_xoxo ???

OpenStudy (anonymous):

no :(

OpenStudy (ash2326):

Which part you have doubt?

OpenStudy (anonymous):

oh, nvm i get it now :)

OpenStudy (anonymous):

thanks!

OpenStudy (ash2326):

Cool